### v1

@@ 1 Introduction
Conventional wisdom suggests that the interaction of people with different backgrounds and perspectives will often lead to socially desirable outcomes. Indeed, economists have advanced intuitive arguments linking diversity to successful problem-solving and improved decision-making, relying on the idea that different perspectives and capabilities naturally serve as complements, enriching and refining each other (e.g., Hong2001; Hong2004; Page). There is much less focus on the productivity implications of open and persistent disagreement, characterized by strongly conflicting views and interpretations of a problem. In this paper we address the following question: can disagreement in a team of innovators boost the team’s output? When should we expect this to happen? We bring to light a force specific to disagreement: the incentive to persuade others through production breakthroughs. In short, disagreement increases output because optimistic agents work harder initially to convince pessimists, and this incentive outweighs the reduced effort of pessimists.
@@
Seminal contributions to the economic literature have warned us about the perils of preference disagreement, shown to create impasse and inefficiencies in many domains of social interaction. [^1] For instance, disagreement has been shown to impede decision-making and compromise economic outcomes in social choice (e.g. Arrow1951), communication (e.g., CS), public finance and public good provision (e.g, AlesinaT1990; AlesinaLaF2005). Yet, history suggests a link between the conflict of worldviews and greater innovation. The development of the first iPhone was, anecdotally, a story of disagreement (iPhone; Grant2021). Steve Jobs initially thought the product would only appeal to a “pocket protector” crowd, and saw the project as a dead end. A team of hard-working engineers — with Apple’s design chief Jonathan Ive on their side — had a very different opinion. They believed that the touchscreen technology would represent a paradigm shift for the industry. Job’s skepticism meant that the team needed to design a prototype so good that it would have been impossible for him not to change his mind. [^2] According to Grant2021: “Fadell and his engineers chipped away at the resistance by building early prototypes in secret, showing Jobs demos, and refining their designs.” Some argue that Job’s open disagreement culture was the key to the company’s success in those years (e.g., Scott2017; Grant2021).
@@
Beyond the iPhone anecdote, the power of disagreement has been recognized in many innovation-related contexts. Open disagreement between scientists has led to the production of more and better theories, resulting in the belief that scientific skepticism — the tendency to challenge and falsify existing theories — lies naturally at the heart of scientific progress (Kuhn1962). Peers’ skepticism has typically motivated philosophers to design sophisticated arguments in favor of their worldview, to convince others to adopt it. [^3] For instance, determined to convince skeptics about the existence of God, Anselm of Canterbury designed the first ontological argument, a class of arguments that has fascinated philosophers for almost a thousand years (see https://plato.stanford.edu/entries/ontological-arguments/). Artists have often found in partners’ disagreement and competition of ideas a motivating force inspiring them to innovate and often reach success. Metallica illustrate this point using the case study of the heavy metal band Metallica and many of their historical producers (most notably, producer Bob Rock): “Both sides [Metallica and Rock] wanted to make the best record in the world but disagreed on what it should be like and how it should be done. […] the most constructive outcomes of conflicts realized when Metallica collaborated with its partners but at the same time also competed with them. The partners first competed to find the best idea, after which the best idea was further developed together.”
@@
Outside the realm of intellectual and artistic debates, the space race was perhaps one of the starkest examples of how the collision of conflicting beliefs about the ideal economy and society spurred technological investment, leading to remarkable innovations in an attempt to prove the superiority of one worldview over the other (TechPolitik). [^4] According to Brian C. Odom, Nasa Chief Historian: “In the global South, you had a lot of countries becoming independent from former colonial powers. What system would they follow? Would they follow the U.S. liberal democracy or would they follow the Soviet example of communism? Kennedy saw the race to the moon as a way to demonstrate American technological power and the benefit of one system over another.” See https://www.space.com/space-race.html. The more recent standards wars in the Tech industry — battles for market dominance between incompatible technologies — can be thought of in a similar fashion. [^5] R&D breakthroughs are recognized as typical ways of persuading the key players and winning such wars. This was the case, for instance, in the battle between the AC and DC technologies for the generation and distribution of power. See ShapiroVarian for a discussion of standard wars.
@@
This paper explores the productivity implications of disagreement in a simple class of games sharing the above three characteristics. A team of coworkers engage in a two-period production game. In each period, they simultaneously and independently choose levels and allocation of costly effort across one or more production technologies. At the end of each period, the team’s output is realized, and payoffs — increasing in such output — accrue. We make the following key assumptions. First, coworkers are uncertain about the returns on effort of the available production technologies — high or low — and can start the game with different prior beliefs about these returns. Team members are aware of each others’ priors, and whenever these priors differ, they think others are wrong: they disagree. [^6] Note that awareness of pior disagreement does not contradict Aumann’s “agreeing to disagree” result, which assumes a common prior. Second, the higher the effort invested in a technology, the more likely the technology is to produce a return (breakthrough) and reveal its value. Third, players observe past breakthroughs and update their beliefs accordingly.
@@ 1.1 Literature Review
At a high level, our paper is related to the economic literature on team diversity (e.g., Hong2001; Hong2004; Page; Dong2022). We highlight how the interaction between coworkers’ strategic incentives and their perspectives on the production process can define whether team members with different views will, on average, outperform homogeneous teams. In contrast with a large part of the diversity literature, the driving force of our results is not just team heterogeneity, but players’ awareness about disagreement: the benefits from disagreement highlighted in this paper vanish if each player behaved under the (wrong) assumption that others shared their prior belief. Hence, we highlight the importance of open disagreement.
Some of our ideas are related to the literature on Bayesian learning with different priors (e.g., Hart2020; Kartik2021), and its applications to principal-agent problems of evidence collection (e.g., Che2009; VanDerSteen) and delegation (e.g., Hirsch2016). In particular, Che2009 and VanDerSteen show that disagreement with the principal can motivate an expert with aligned preferences to collect more evidence. [^7] VanDerSteen also finds that disagreement between a manager and an employee tends to reduce delegation, hinder motivation, and lower satisfaction. By focusing on different strategic incentives, we provide a complementary, more optimistic picture of the effects of disagreement on team effort and production. We complement this literature with an analysis of team formation and teamwork, where agents persuade each other through successful production histories. Importantly, we show why an output-maximizing manager could benefit from forming disagreeing teams even when she is perfectly informed about the state and therefore does not value additional information. Additionally, our analysis suggests that the persuasion incentives of disagreeing coworkers might reduce the average output of skeptical team members.
Our analysis shares qualitative similarities with the literature on exponential bandits (e.g., Keller2005; Moroni21), and in particular Dong2018. By assuming the unobservability of effort, we shut down signaling, the driving mechanism of Dong2018, isolate the effect of persuasion through output, absent in their paper, and reach different conclusions. [^8] Dong2018 finds that asymmetric information about the risky arm can increase experimentation in a team. Her results are driven by informational asymmetries and signaling instead of open disagreement and persuasion. In Dong2018, the player with bad news increases effort in order to pool with the player with good news. In our model, the optimistic player takes the lead, working harder to bring the skeptic on board.
@@ 2.1 The Production Game
A team of coworkers engages in productive activity over two periods. In each period $t=1,2$, each team member i=1,,Ni=1,...,N chooses how much effort $e^{i}_{t}\in[0,1]$ to allocate to production. The cost of effort is quadratic [^9] The functional form assumption comes with the benefit of nice closed-form solutions for equilibrium effort levels, and allows us to better isolate the productivity distortions due to disagreement. Our main results generalize to convex cost functions with weakly negative third derivative. and described by $c(e^{i}_{t})=\frac{1}{2}(e_{t}^{i})^{2}$ and, for any level of aggregate effort $e_{t}=\sum_{i=1}^{N}e^{i}_{t}$ exerted by the team in period $t$, the productive activity yields a return $\omega\in\{R,r\}$ with probability $\min\{e_{t},1\}$ and no return with the remaining probability. [^10] The assumption that the probability of a return increases linearly in the aggregate team effort shuts down effort complementarities due to the production technology, and allows us to isolate the information-driven implications of disagreement that are the focus of this paper. See Prat2002 for a study of how team heterogeneity affects productivity in the presence of technology-induced effort complementarities. We denote by $y_{t}\in\{R,r,0\}$ the value of team production in period $t$, where $y_{t}=0$ means that the team effort was not successful. We maintain the assumption that, conditional on effort levels, the arrival of returns is independent across periods.
The type of return $\omega$ of the production technology is fixed throughout the game and initially uncertain. When effort is successful, high-return technologies yield $\omega=R$, while low-return technologies yield $\omega=r$, with $R>r>0$. Player $i$ attaches prior probability $p^{i}\in[0,1]$ to $\omega=R$, and players can start the game with different priors. Dispensing with the common prior hypothesis allows us to study how the game outcomes depend on initial team disagreement. Furthermore, we assume that players know each other’s priors: whenever there is disagreement in the team, it is “open disagreement.” Note that this awareness of prior disagreement is not inconsistent with Aumann’s impossibility of “agreeing to disagree,” as such a result holds for posterior beliefs when players hold a common prior.
@@
Let $\pi^{i}$ denote the probability that $\omega=R$ according to the belief of player i=1,,Ni=1,...,N at the beginning of $t=2$. If $y_{1}>0$ then $\pi^{i}=\mathbbm{1}(y_{1}=R)$. If $y_{1}=0$ then $\pi^{i}=p^{i}$.
The assumption requires that, after observing a high (low) return, each player assigns probability $1$ to the technology being of the high-return (low-return) type and that, in the absence of a return, players do not update their prior. This requirement allows for Bayesian players, as well as players following other updating rules, including paradigm-shift rules (Ortoleva2012; Galperti2019).
@@ 3 The Disagreement Dividend
From the perspective of player $i$, learning has two implications: (i) initial disagreement is resolved, and (ii) the prior belief $p^{i}$ is updated, leading to a better allocation of effort. Equation 3 decomposes the effect of learning in these two components. Resolving initial disagreement means that $j^{\prime}s$ belief is corrected from $p^{j}$ to $p^{i}$, yielding a change in her future effort from $e^{\star}_{2}(p^{j})$ to $e^{\star}_{2}(p^{i})$. With this belief correction, the present value of $i$’s expected second-period payoff changes by the second term on the right-end side of equation 3. Note that $e^{\star}_{2}(p^{i})-e^{\star}_{2}(p^{j})$ is $i$’s expectation about $j$’s effort change from learning: since $\mathbb{E}_{p^{i}}\left[\pi^{j}|y_{1}>0\right]=p^{i}$, it has
@@
This highlights the key persuasion mechanism in place in a disagreeing team. When Ann is more optimistic than Bob, she expects a return arrival, on average, to confirm her belief and make Bob more optimistic. Optimism motivates Bob to work harder in the future. As Ann benefits from Bob’s future hard work, she works harder early on to increase the likelihood of a return in the first period, compared to the case when Bob is like-minded. The reverse implication works for Bob: as he expects a return to make Ann more pessimistic, he has the incentive to work less in the first period than when paired with a like-minded coworker.
@@ 4 Team Formation
What gives rise to this anti-monotone team formation plan? In the appendix we show that
@@ 5 A Tale of Two Methods
The intuition of the proposition is the following. In stark contrast with the result of proposition 1, when there is competition of ideas, pessimism about the technology favored by the coworker does not discourage effort, but motivates a team member to work harder. In particular, we show that at $t=1$, each worker always operates the technology she’s optimistic about, anticipating that an early success will persuade her teammate to switch to her “superior” production approach in the future. To provide a more formal intuition, the first order condition for the first-period equilibrium effort level of $i=1,2$ when $p^{\star}=(1,1)$ becomes
@@
The takeaway of this section is simple and yet, we believe, important and non-trivial. When two similarly good production technologies are available, disagreement on the best way to produce might increase group production. We focused on a mechanism that builds on the idea of different perspectives and shows that these differences can be useful even when they do not complement each other, that is, when there is competition of ideas. When they have the common goal of increasing each others’ output, disagreeing people will challenge each other, and work harder to prove, with their successes, that their perspective is valid, beneficial, and worth adopting. The resulting increase in group output can make all team members better off.
@@ 6 Conclusion
We have shown that disagreement within a group of economic agents who repeatedly engage in a productive activity can increase the aggregate output of the group. This relation between disagreement and productivity helps us benchmark our results with the main findings of the theoretical literature on diversity and problem solving (e.g., Hong2001; Hong2004). In this literature, different perspectives are seen as an asset, but under two assumptions. First, team members must be able to cooperate and combine their perspectives in a productive way. Second, and related, different perspectives must not lead to different goals.

### v2

@@ 1 Introduction
Conventional wisdom suggests that the interaction of people with different backgrounds and perspectives will often lead to socially desirable outcomes. Indeed, economists have advanced intuitive arguments linking diversity to successful problem-solving and improved decision-making, relying on the idea that different perspectives and capabilities naturally serve as complements, enriching and refining each other (e.g., Hong and Page, 2001, 2004; Page, 2007). There is much less focus on the productivity implications of open and persistent disagreement, characterized by strongly conflicting views and interpretations of a problem. In this paper we address the following question: can disagreement in a team of innovators boost the team’s output? When should we expect this to happen? We bring to light a force specific to disagreement: the incentive to persuade others through production breakthroughs. In short, disagreement increases output because optimistic agents work harder initially to convince pessimists, and this incentive outweighs the reduced effort of pessimists.
@@
Seminal contributions to the economic literature have warned us about the perils of preference disagreement, shown to create impasse and inefficiencies in many domains of social interaction. [^1] For instance, disagreement has been shown to impede decision-making and compromise economic outcomes in social choice (e.g. Arrow, 1951), communication (e.g., Crawford and Sobel, 1982), public finance and public good provision (e.g, Alesina and Tabellini, 1990; Alesina and La Ferrara, 2005). Yet, history suggests a link between the conflict of worldviews and greater innovation. The development of the first iPhone was, anecdotally, a story of disagreement (Merchant, 2018; Grant, 2021). Steve Jobs initially thought the product would only appeal to a “pocket protector” crowd, and saw the project as a dead end. A team of hard-working engineers — with Apple’s design chief Jonathan Ive on their side — had a very different opinion. They believed that the touchscreen technology would represent a paradigm shift for the industry. Job’s skepticism meant that the team needed to design a prototype so good that it would have been impossible for him not to change his mind. [^2] According to Grant (2021): “Fadell and his engineers chipped away at the resistance by building early prototypes in secret, showing Jobs demos, and refining their designs.” Some argue that Job’s open disagreement culture was the key to the company’s success in those years (e.g., Scott, 2017; Grant, 2021).
@@
Beyond the iPhone anecdote, the power of disagreement has been recognized in many innovation-related contexts. Open disagreement between scientists has led to the production of more and better theories, resulting in the belief that scientific skepticism — the tendency to challenge and falsify existing theories — lies naturally at the heart of scientific progress (Kuhn, 1962). Peers’ skepticism has typically motivated philosophers to design sophisticated arguments in favor of their worldview, to convince others to adopt it. [^3] For instance, determined to convince skeptics about the existence of God, Anselm of Canterbury designed the first ontological argument, a class of arguments that has fascinated philosophers for almost a thousand years (see https://plato.stanford.edu/entries/ontological-arguments/). Artists have often found in partners’ disagreement and competition of ideas a motivating force inspiring them to innovate and often reach success. Salmela and Oikkonen (2022) illustrate this point using the case study of the heavy metal band Metallica and many of their historical producers (most notably, producer Bob Rock): “Both sides [Metallica and Rock] wanted to make the best record in the world but disagreed on what it should be like and how it should be done. […] the most constructive outcomes of conflicts realized when Metallica collaborated with its partners but at the same time also competed with them. The partners first competed to find the best idea, after which the best idea was further developed together.”
@@
Outside the realm of intellectual and artistic debates, the space race was perhaps one of the starkest examples of how the collision of conflicting beliefs about the ideal economy and society spurred technological investment, leading to remarkable innovations in an attempt to prove the superiority of one worldview over the other (Center and Bates, 2009). [^4] According to Brian C. Odom, Nasa Chief Historian: “In the global South, you had a lot of countries becoming independent from former colonial powers. What system would they follow? Would they follow the U.S. liberal democracy or would they follow the Soviet example of communism? Kennedy saw the race to the moon as a way to demonstrate American technological power and the benefit of one system over another.” See https://www.space.com/space-race.html. The more recent standards wars in the Tech industry — battles for market dominance between incompatible technologies — can be thought of in a similar fashion. [^5] R&D breakthroughs are recognized as typical ways of persuading the key players and winning such wars. This was the case, for instance, in the battle between the AC and DC technologies for the generation and distribution of power. See Shapiro and Varian (1999) for a discussion of standard wars.
@@
This paper explores the productivity implications of disagreement in a simple class of games sharing the above three characteristics. A team of coworkers engage in a two-period production game. In each period, they simultaneously and independently choose levels and allocation of costly effort across one or more production technologies. At the end of each period, the team’s output is realized, and payoffs — increasing in such output — accrue. We make the following key assumptions. First, coworkers are uncertain about the returns on effort of the available production technologies — high or low — and can start the game with different prior beliefs about these returns. Team members are aware of each others’ priors, and whenever these priors differ, they think others are wrong: they disagree. [^6] Note that awareness of pior disagreement does not contradict Aumann (1976)’s “agreeing to disagree” result, which assumes a common prior. Second, the higher the effort invested in a technology, the more likely the technology is to produce a return (breakthrough) and reveal its value. Third, players observe past breakthroughs and update their beliefs accordingly.
@@ 1.1 Literature Review
At a high level, our paper is related to the economic literature on team diversity (e.g., Hong and Page, 2001, 2004; Page, 2007; Dong and Mayskaya, 2022). We highlight how the interaction between coworkers’ strategic incentives and their perspectives on the production process can define whether team members with different views will, on average, outperform homogeneous teams. In contrast with a large part of the diversity literature, the driving force of our results is not just team heterogeneity, but players’ awareness about disagreement: the benefits from disagreement highlighted in this paper vanish if each player behaved under the (wrong) assumption that others shared their prior belief. Hence, we highlight the importance of open disagreement.
Some of our ideas are related to the literature on Bayesian learning with different priors (e.g., Hart and Rinott, 2020; Kartik et al., 2021), and its applications to principal-agent problems of evidence collection (e.g., Che and Kartik, 2009; Van den Steen, 2010) and delegation (e.g., Hirsch, 2016). In particular, Che and Kartik (2009) and Van den Steen (2010) show that disagreement with the principal can motivate an expert with aligned preferences to collect more evidence. [^7] Van den Steen (2010) also finds that disagreement between a manager and an employee tends to reduce delegation, hinder motivation, and lower satisfaction. By focusing on different strategic incentives, we provide a complementary, more optimistic picture of the effects of disagreement on team effort and production. We complement this literature with an analysis of team formation and teamwork, where agents persuade each other through successful production histories. Importantly, we show why an output-maximizing manager could benefit from forming disagreeing teams even when she is perfectly informed about the state and therefore does not value additional information. Additionally, our analysis suggests that the persuasion incentives of disagreeing coworkers might reduce the average output of skeptical team members.
Our analysis shares qualitative similarities with the literature on exponential bandits (e.g., Keller et al., 2005; Moroni, 2022), and in particular Dong (2018). By assuming the unobservability of effort, we shut down signaling, the driving mechanism of Dong (2018), isolate the effect of persuasion through output, absent in their paper, and reach different conclusions. [^8] Dong (2018) finds that asymmetric information about the risky arm can increase experimentation in a team. Her results are driven by informational asymmetries and signaling instead of open disagreement and persuasion. In Dong (2018), the player with bad news increases effort in order to pool with the player with good news. In our model, the optimistic player takes the lead, working harder to bring the skeptic on board.
@@ 2.1 The Production Game
A team of coworkers engages in productive activity over two periods. In each period $t=1,2$, each team member i=1,,Ni=1,...,N chooses how much effort $e^{i}_{t}\in[0,1]$ to allocate to production. The cost of effort is quadratic [^9] The functional form assumption comes with the benefit of nice closed-form solutions for equilibrium effort levels, and allows us to better isolate the productivity distortions due to disagreement. Our main results generalize to convex cost functions with weakly negative third derivative. and described by $c(e^{i}_{t})=\frac{1}{2}(e_{t}^{i})^{2}$ and, for any level of aggregate effort $e_{t}=\sum_{i=1}^{N}e^{i}_{t}$ exerted by the team in period $t$, the productive activity yields a return $\omega\in\{R,r\}$ with probability $\min\{e_{t},1\}$ and no return with the remaining probability. [^10] The assumption that the probability of a return increases linearly in the aggregate team effort shuts down effort complementarities due to the production technology, and allows us to isolate the information-driven implications of disagreement that are the focus of this paper. See Prat (2002) for a study of how team heterogeneity affects productivity in the presence of technology-induced effort complementarities. We denote by $y_{t}\in\{R,r,0\}$ the value of team production in period $t$, where $y_{t}=0$ means that the team effort was not successful. We maintain the assumption that, conditional on effort levels, the arrival of returns is independent across periods.
The type of return $\omega$ of the production technology is fixed throughout the game and initially uncertain. When effort is successful, high-return technologies yield $\omega=R$, while low-return technologies yield $\omega=r$, with $R>r>0$. Player $i$ attaches prior probability $p^{i}\in[0,1]$ to $\omega=R$, and players can start the game with different priors. Dispensing with the common prior hypothesis allows us to study how the game outcomes depend on initial team disagreement. Furthermore, we assume that players know each other’s priors: whenever there is disagreement in the team, it is “open disagreement.” Note that this awareness of prior disagreement is not inconsistent with Aumann (1976)’s impossibility of “agreeing to disagree,” as such a result holds for posterior beliefs when players hold a common prior.
@@
Let $\pi^{i}$ denote the probability that $\omega=R$ according to the belief of player i=1,,Ni=1,...,N at the beginning of $t=2$. If $y_{1}>0$ then $\pi^{i}=\mathbbm{1}(y_{1}=R)$. If $y_{1}=0$ then $\pi^{i}=p^{i}$, where $p^{i}$ is the probability that $\omega=R$ according to the prior of player $i$.
The assumption requires that, after observing a high (low) return, each player assigns probability $1$ to the technology being of the high-return (low-return) type and that, in the absence of a return, players do not update their prior. This requirement allows for Bayesian players, as well as players following other updating rules, including paradigm-shift rules (Ortoleva, 2012; Galperti, 2019).
@@ 3 The Disagreement Dividend
From the perspective of player $i$, learning has two implications: (i) initial disagreement is resolved, and (ii) the prior belief $p^{i}$ is updated, leading to a better allocation of effort. Equation 3 decomposes the effect of learning in these two components. Resolving initial disagreement means that $j^{\prime}s$ belief is corrected from $p^{j}$ to $p^{i}$, yielding a change in her future effort from $e^{\star}_{2}(p^{j})$ to $e^{\star}_{2}(p^{i})$. With this belief correction, the present value of $i$’s expected second-period payoff changes by the second term on the right-hand side of equation 3. Note that $e^{\star}_{2}(p^{i})-e^{\star}_{2}(p^{j})$ is $i$’s expectation about $j$’s effort change from learning: since $\mathbb{E}_{p^{i}}\left[\pi^{j}|y_{1}>0\right]=p^{i}$, it has
@@
This highlights the key persuasion mechanism in place in a disagreeing team. When Ann is more optimistic than Bob, she expects a return arrival, on average, to confirm her belief and make Bob more optimistic. Such an increase in optimism motivates Bob to work harder in the future, and Ann benefits from Bob’s future hard work. Hence, Bob’s pessimism motivates Ann to work harder early on to increase the likelihood of a return in the first period, compared to the case when Bob is like-minded. The reverse implication works for Bob: as he expects a return to make Ann more pessimistic, he has the incentive to work less in the first period than when paired with a like-minded coworker.
@@ 4 Team Formation
What gives rise to this negative sorting? In the appendix we show that
@@ 5 A Tale of Two Methods
The intuition of the proposition is the following. In stark contrast with the result of proposition 1, when there is competition of ideas, pessimism about the technology favored by the coworker does not discourage effort, but motivates a team member to work harder. In particular, we show that at $t=1$, each worker always operates the technology she’s optimistic about, anticipating that an early success will persuade her teammate to switch to her “superior” production approach in the future. More formally, the first order condition for the first-period equilibrium effort level of $i=1,2$ becomes
@@
The takeaway of this section is simple and yet, we believe, important and non-trivial. When two similarly good production technologies are available, disagreement on the best way to produce might increase group production. We focused on a mechanism that builds on the idea of different perspectives and shows that these differences can be useful even when they do not complement each other, that is, when there is competition of ideas. When they have the common goal of increasing each other’s productivity, disagreeing people will challenge each other, and work harder to prove, with their successes, that their own perspective is valid, beneficial, and worth adopting. The resulting increase in group output can make all team members better off.
@@ 6 Conclusion
We have shown that disagreement within a group of economic agents who repeatedly engage in a productive activity can increase the aggregate output of the group. This relation between disagreement and productivity helps us benchmark our results with the main findings of the theoretical literature on diversity and problem solving (e.g., Hong and Page, 2001, 2004). In this literature, different perspectives are seen as an asset, but under two assumptions. First, team members must be able to cooperate and combine their perspectives in a productive way. Second, and related, different perspectives must not lead to different goals.